OpenStudy (anonymous):

* Advance Algebra With Financial Applications * Fiona is purchasing a condominium and is financing $305,000 with a 30-year 5/1 ARM at 4.65% with a 1/12 cap structure. What will her payments be at the beginning of year 6? $1436.86 $1572.69 $1608.51 $1736.26

OpenStudy (anonymous):

I know I need to use this M=B[((i)(1+i)^(n*t))/((1+i)^(n*t)-(1))]

OpenStudy (anonymous):

I just wanna know what's going to be "i" ?

OpenStudy (amistre64):

for 5 years the i is set; then for each year after it goes up by 1 to a cap of at most 12

OpenStudy (amistre64):

my process is to determine the payments for the first 5 years then balance out the loan after 5 years of those payments then use the remaining balance to recalulate the payments in year 6 at +1 interst

OpenStudy (amistre64):

$305,000 with a 30-year 5/1 ARM at 4.65% P = 305500(k^12*30)(1-k)/(1-k^(12*30)), k=1+.0465/12 Balance after 5 years is: B5 = 305500k^(5*12) - P(1-k^(5*12))/(1-k). k=1+.0456/12

OpenStudy (amistre64):

use B5, and the remaining time (25 years) and an interest of 5.65 to recalculate the payment

OpenStudy (amistre64):

P = 1575.27 for the first 5 years 5 year balance gets us to: 277 634 Po = 277634k^(12*25)(1-k)/(1-k^(12*25)), k=1+.0565/12, or about 1729.88

OpenStudy (amistre64):

due to some rnding errors ... id say thats close enough :)

OpenStudy (amistre64):

1736 without the rnds

OpenStudy (anonymous):

Thank you soo much !

OpenStudy (amistre64):

yep